OpenStudy (anonymous):

You give your dad $500 at the end of every year for the next 10 years to save for you. At the end of the 10 years, he gives you $8,000! What rate of interest did your dad pay you?

OpenStudy (johnweldon1993):

This sounds like a simple interest problem I = PRT 'I' = interest earned = what father gave you - what you gave him 'P' = principal rate = original rate = 500 'R' = rate (what we want to solve for) 'T' = time ( in years) We want to solve for R so...rearrange this for R \[R = \frac{ I }{ PT }\] We know I = 8000 - (500 * 10) = 3000 T = 10 years P = 500 dollars so \[R = \frac{ 3000 }{ 500 * 10 }\] What does R = ...?

OpenStudy (anonymous):

R=.6, So it would be at a 60% interest rate?

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

thank you!

OpenStudy (johnweldon1993):

No problem!